### v1

@@ 3 Nonlinear iteration scheme
$\bm{r}(\bar{\bm{u}}):=\bm{b}-{\cal A}(\bar{\bm{u}})\equiv\begin{bmatrix}\bar{\bm{u}}^{0}\\ \bm{0}\\ \vdots\\ \bm{0}\end{bmatrix}-\begin{bmatrix}&I&\\ &-\Phi(\,\cdot\,)&I\\ &&\ddots&\ddots\\ &&&-\Phi(\,\cdot\,)&I\end{bmatrix}\begin{bmatrix}\bar{\bm{u}}^{0}\\ \bar{\bm{u}}^{1}\\ \vdots\\ \bar{\bm{u}}^{n_{t}-1}\end{bmatrix}=\bm{0}.$ (27)
Here $\bar{\bm{u}}=\big{(}\bar{\bm{u}}^{0},\bar{\bm{u}}^{1},\ldots,\bar{\bm{u}}^{n_{t}-1}\big{)}^{\top}\in\mathbb{R}^{n_{x}n_{t}}$ is created from concatenating the cell-averages we seek across the time domain, ${\cal A}\colon\mathbb{R}^{n_{x}n_{t}}\to\mathbb{R}^{n_{x}n_{t}}$ is the nonlinear space-time discretization operator we seek to invert, and $\bm{b}\in\mathbb{R}^{n_{x}n_{t}}$ contains the initial data for the problem. For some approximate solution $\bar{\bm{u}}_{k}\approx\bar{\bm{u}}$, with iteration index $k$, the residual (27) is denoted as $\bm{r}_{k}:=\bm{r}(\bar{\bm{u}}_{k})$.
In this paper, we iteratively solve (27) using the linearly preconditioned residual correction scheme outlined in Algorithm 1. The remainder of this section discusses in greater detail the development of Algorithm 1; however, we first use it to contextualize the key contributions of this paper. In Algorithm 1, $P_{k}$ is a certain linearization of the space-time discretization ${\cal A}$ in (27), and the development of this linearization in Section 4 is the first key contribution of this work; specifically, $P_{k}$ is realized as a certain non-standard discretization of the linear conservation law (cons-lin). In Algorithm 1 linear systems of the form $P_{k}\bar{\bm{e}}_{k}^{\textrm{lin}}=\bm{r}_{k}$ are solved either directly with sequential time-stepping or approximately with MGRIT. In this work we focus on the latter option since we are concerned with parallel-in-time solvers. The development of an effective MGRIT solver for $P_{k}$ is the second key contribution of this work, and is the subject of Section 5. Our MGRIT solver is based on a novel, modified FV semi-Lagrangian coarse-grid discretization of (cons-lin) which is an extension of our modified finite-difference (FD) semi-Lagrangian discretizations in [14, 16, 17] for non-conservative linear hyperbolic PDEs. The efficacy of our solution methodology is demonstrated in Section 6 by way of numerical tests on challenging nonlinear problems containing shock and rarefaction waves.
@@
5:         ${\bar{\bm{e}}}_{k}^{\textrm{lin}}\leftarrow P_{k}^{-1}{\bm{r}}_{k}$triangleright Direct solve (sequential in time)
@@
7:         ${\bar{\bm{e}}}_{k}^{\textrm{lin}}\leftarrow\textrm{MGRIT}(P_{k},{\bm{r}}_{k})$triangleright Approx. solve (parallel in time)
@@
Since ${\cal A}$ in (27), and consequently $\nabla_{\bar{\bm{u}}}{\cal A}(\bar{\bm{u}})$, corresponds to a one-step discretization in time, we impose that $P_{k}$ has the corresponding sparsity structure:
$P_{k}=\begin{bmatrix}&I\\ &-\Phi^{\textrm{lin}}(\bar{\bm{u}}_{k}^{0})&I\\ &&\ddots&\ddots\\ &&&-\Phi^{\textrm{lin}}(\bar{\bm{u}}_{k}^{n_{t}-2})&I\end{bmatrix}\in\mathbb{R}^{n_{x}n_{t}\times n_{x}n_{t}}.$ (33)
@@
$\Phi^{\textrm{lin}}(\bar{\bm{u}}^{n}_{k})=\nabla_{\bm{\alpha}}\widetilde{\Phi}(\bm{\alpha},\bm{\beta})\big{|}_{(\bm{\alpha},\bm{\beta})=(\bar{\bm{u}}^{n}_{k},\bar{\bm{u}}^{n}_{k})}.$ (34)
In the coming sections we detail specifically how $\widetilde{\Phi}$ is chosen (see e.g., (43)). Moving forward, to maintain a uniform and simple notation, we continue to refer to the linearized time-stepping operator $\Phi^{\textrm{lin}}(\bar{\bm{u}}^{n}_{k})$ in (34) as the “Jacobian of $\Phi$” and use the corresponding notation “$\nabla\Phi$,” also when the Jacobian does not exist. In cases where it does not exist, “Jacobian of $\Phi$” and “$\nabla\Phi$” are to be understood in the sense described above and as written in (34).
Finally, our numerical tests indicate that the linearized iteration (28) may benefit from the addition of nonlinear block relaxations of the type that would be done on (27) if the system were to be solved with FAS MGRIT. Hence we add Line 2 in Algorithm 1. To this end, let $m\in\mathbb{N}\setminus\{1\}$ induce a coarse-fine splitting (CF-splitting) of the time points such that every $m$th time point is a C-point, and all other time points are F-points. Then, nonlinear C- and F-relaxations update the current approximation $\bar{\bm{u}}_{k}$ such that the nonlinear residual is zero at C- and F-points, respectively [37]. A C-relaxation requires time-stepping to each C-point from its preceding F-point, and an F-relaxation requires time-stepping from each C-point to its succeeding $m-1$ F-points. In practice we only do an F-relaxation in Line 2, but note that any stronger relaxation ending in F is also possible.
@@ 4.1 Linearization I: Time integration
$\big{[}\nabla_{\bar{\bm{u}}^{n}}\Phi(\bar{\bm{u}}^{n})\big{]}\bar{\bm{e}}^{n}=\big{[}\nabla_{\bar{\bm{u}}^{n}}F(\bar{\bm{u}}^{n})\big{]}\bar{\bm{e}}^{n}=\underbrace{\big{[}I+\delta t\nabla_{\bar{\bm{u}}^{n}}L(\bar{\bm{u}}^{n})\big{]}}_{=:F^{\textrm{lin}}(\bar{\bm{u}}^{n})}\bar{\bm{e}}^{n}.$ (35)
@@
$\nabla_{\bar{\bm{u}}^{n}}\Phi(\bar{\bm{u}}^{n})\bar{\bm{e}}^{n}=\frac{1}{3}\bar{\bm{e}}^{n}+\frac{2}{3}\nabla_{\bar{\bm{u}}^{n}}F\big{(}\bar{\bm{u}}^{n,2}\big{)}\bar{\bm{e}}^{n}=\frac{1}{3}\bar{\bm{e}}^{n}+\frac{2}{3}F^{\textrm{lin}}\big{(}\bar{\bm{u}}^{n,2}\big{)}\underbrace{\nabla_{\bar{\bm{u}}^{n}}\bar{\bm{u}}^{n,2}\bar{\bm{e}}^{n}}_{=:\bar{\bm{e}}^{n,2}}.$ (36)
@@
$\bar{\bm{e}}^{n,2}=\frac{3}{4}\bar{\bm{e}}^{n}+\frac{1}{4}F^{\textrm{lin}}\big{(}\bar{\bm{u}}^{n,1}\big{)}\underbrace{\nabla_{\bar{\bm{u}}^{n}}\bar{\bm{u}}^{n,1}\bar{\bm{e}}^{n}}_{=:\bar{\bm{e}}^{n,1}},\quad\textrm{and}\quad\bar{\bm{e}}^{n,1}=F^{\textrm{lin}}(\bar{\bm{u}}^{n})\bar{\bm{e}}^{n}.$ (37)
@@ 4.2 Linearization II: Numerical flux
$-\nabla_{\bar{\bm{u}}}\big{(}L(\bar{\bm{u}})\big{)}_{i}\bar{\bm{e}}=\frac{\nabla_{\bar{\bm{u}}}\widehat{f}_{i+1/2}(\bar{\bm{u}})\bar{\bm{e}}-\nabla_{\bar{\bm{u}}}\widehat{f}_{i-1/2}(\bar{\bm{u}})\bar{\bm{e}}}{h}.$ (38)
@@
$\nabla_{\bar{\bm{u}}}\widehat{f}_{i+1/2}^{\rm(global)}(\bar{\bm{u}})\bar{\bm{e}}=\nabla_{\bar{\bm{u}}}\widehat{f}^{\rm(global)}\big{(}u_{i+1/2}^{-}(\bar{\bm{u}}),u_{i+1/2}^{+}(\bar{\bm{u}})\big{)}\bar{\bm{e}},$ (39)
$=\left(\frac{\partial\widehat{f}^{\rm(global)}}{\partial u_{i+1/2}^{-}}\nabla_{\bar{\bm{u}}}u_{i+1/2}^{-}(\bar{\bm{u}})+\frac{\partial\widehat{f}^{\rm(global)}}{\partial u_{i+1/2}^{+}}\nabla_{\bar{\bm{u}}}u_{i+1/2}^{+}(\bar{\bm{u}})\right)\bar{\bm{e}},$ (40)
$\begin{split}&=\frac{1}{2}\left(f^{\prime}\big{(}u_{i+1/2}^{-}\big{)}+\nu^{\rm(global)}\right)\Big{(}\nabla_{\bar{\bm{u}}}u_{i+1/2}^{-}(\bar{\bm{u}})\bar{\bm{e}}\Big{)}\\ &\hskip 34.44434pt+\frac{1}{2}\left(f^{\prime}\big{(}u_{i+1/2}^{+}\big{)}-\nu^{\rm(global)}\right)\Big{(}\nabla_{\bar{\bm{u}}}u_{i+1/2}^{+}(\bar{\bm{u}})\bar{\bm{e}}\Big{)}\end{split},$ (41)
$\begin{split}&=\widehat{f}^{\textrm{lin}}\Big{(}f^{\prime}\big{(}u_{i+1/2}^{-}\big{)},\nabla_{\bar{\bm{u}}}u_{i+1/2}^{-}(\bar{\bm{u}})\bar{\bm{e}},f^{\prime}\big{(}u_{i+1/2}^{+}\big{)},\nabla_{\bar{\bm{u}}}u_{i+1/2}^{+}(\bar{\bm{u}})\bar{\bm{e}},\nu^{\rm(global)}\Big{)}.\end{split}$ (42)
@@
$\begin{split}\widetilde{f}_{i+1/2}^{\rm(local)}(\bm{\alpha},\bm{\beta})&:=\frac{1}{2}\Big{[}f(u_{i+1/2}^{-}(\bm{\alpha}))+f(u_{i+1/2}^{+}(\bm{\alpha}))\Big{]}\\ &\hskip 34.44434pt+\frac{1}{2}\nu_{i+1/2}(\bm{\beta})\Big{[}u_{i+1/2}^{-}(\bm{\alpha})-u_{i+1/2}^{+}(\bm{\alpha})\Big{]}\end{split}.$ (43)
@@
$\begin{split}&\nabla_{\bar{\bm{u}}}\widehat{f}_{i+1/2}^{\rm(local)}(\bar{\bm{u}})\bar{\bm{e}}=\nabla_{\bm{\alpha}}\widetilde{f}_{i+1/2}^{\rm(local)}(\bm{\alpha},\bm{\beta})|_{(\bm{\alpha},\bm{\beta})=(\bar{\bm{u}},\bar{\bm{u}})}\bar{\bm{e}}=\\ &\widehat{f}^{\textrm{lin}}\Big{(}f^{\prime}\big{(}u_{i+1/2}^{-}\big{)},\nabla_{\bar{\bm{u}}}u_{i+1/2}^{-}(\bar{\bm{u}})\bar{\bm{e}},f^{\prime}\big{(}u_{i+1/2}^{+}\big{)},\nabla_{\bar{\bm{u}}}u_{i+1/2}^{+}(\bar{\bm{u}})\bar{\bm{e}},\nu_{i+1/2}^{\rm(local)}(\bar{\bm{u}})\Big{)}.\end{split}$ (44)
@@ 4.3 Linearization III: Reconstructions
$\nabla_{\bar{\bm{u}}}u^{-}_{i+1/2}(\bar{\bm{u}})\bar{\bm{e}}$ $=\underbrace{\sum\limits_{\ell=0}^{k-1}b_{i}^{\ell}(\bar{\bm{u}})\big{(}R^{\ell}\bar{\bm{e}}\big{)}_{i}}_{=:R_{i}^{1}}+\underbrace{\sum\limits_{\ell=0}^{k-1}\big{(}R^{\ell}\bar{\bm{u}}\big{)}_{i}\big{(}\nabla_{\bar{\bm{u}}}\,b_{i}^{\ell}(\bar{\bm{u}})\bar{\bm{e}}\big{)}_{i}}_{=:R_{i}^{2}},$ (46)
$\nabla_{\bar{\bm{u}}}u^{+}_{i+1/2}(\bar{\bm{u}})\bar{\bm{e}}$ $=\underbrace{\sum\limits_{\ell=0}^{k-1}\widetilde{b}_{i+1}^{\ell}(\bar{\bm{u}})\big{(}\widetilde{R}^{\ell}\bar{\bm{e}}\big{)}_{i+1}}_{=:\widetilde{R}_{i+1}^{1}}+\underbrace{\sum\limits_{\ell=0}^{k-1}\big{(}\widetilde{R}^{\ell}\bar{\bm{u}}\big{)}_{i+1}\big{(}\nabla_{\bar{\bm{u}}}\,\widetilde{b}_{i+1}^{\ell}(\bar{\bm{u}})\bar{\bm{e}}\big{)}_{i+1}}_{=:\widetilde{R}_{i+1}^{2}}.$ (47)
@@
$k=1:\quad d^{0}=1,\quad\widetilde{d}^{0}=1,$ (48)
$k=2:\quad\big{(}d^{0},d^{1}\big{)}=\Big{(}\frac{2}{3},\frac{1}{3}\Big{)},\quad\big{(}\widetilde{d}^{0},\widetilde{d}^{1}\big{)}=\Big{(}\frac{1}{3},\frac{2}{3}\Big{)}.$ (49)
In this case $R_{i}^{1},\widetilde{R}_{i+1}^{1}$ are simply linear, $(2k-1)$-order reconstructions of $e$. Moreover, the second terms in (46) and (47) vanish, $R_{i}^{2}=\widetilde{R}_{i+1}^{2}=0$, because the weights are independent of $\bar{\bm{u}}$. Recall that these optimal linear weights are used in a 1st-order accurate FV method. However, for higher-order discretizations, these weights are not suitable when the solution lacks regularity because they result in spurious oscillations.
@@
$w^{\ell}_{i}(\bar{\bm{u}})$ $=\frac{\alpha_{i}^{\ell}(\bar{\bm{u}})}{\sum_{\ell=0}^{k-1}\alpha_{i}^{\ell}(\bar{\bm{u}})},\quad\alpha_{i}^{\ell}(\bar{\bm{u}})=\frac{d^{\ell}}{\big{(}\epsilon+\beta_{i}^{\ell}(\bar{\bm{u}})\big{)}^{2}},$ (50)
$\widetilde{w}^{\ell}_{i+1}(\bar{\bm{u}})$ $=\frac{\widetilde{\alpha}_{i+1}^{\ell}(\bar{\bm{u}})}{\sum_{\ell=0}^{k-1}\widetilde{\alpha}_{i+1}^{\ell}(\bar{\bm{u}})},\quad\widetilde{\alpha}_{i+1}^{\ell}(\bar{\bm{u}})=\frac{\widetilde{d}^{\ell}}{\big{(}\epsilon+\beta_{i+1}^{\ell}(\bar{\bm{u}})\big{)}^{2}},$ (51)
with $\epsilon=10^{-6}$. The quantity $\beta^{\ell}$ reflects the smoothness of $u(x)$ over the $\ell$th stencil. For $k=1$, $\beta=0$. For $k=2$ we have [57],
$\beta^{0}_{i}(\bar{\bm{u}})=(\bar{u}_{i+1}-\bar{u}_{i})^{2},\quad\beta^{1}_{i}(\bar{\bm{u}})=(\bar{u}_{i}-\bar{u}_{i-1})^{2}.$ (52)
@@
$\nabla_{\bar{\bm{u}}}u^{\pm}_{i+1/2}(\bar{\bm{u}})\bar{\bm{e}}\approx\frac{u^{\pm}_{i+1/2}(\bar{\bm{u}}+\mu\bar{\bm{e}})-u^{\pm}_{i+1/2}(\bar{\bm{u}})}{\mu},$ (53)
@@
$\nabla_{\bar{\bm{u}}}u^{-}_{i+1/2}(\bar{\bm{u}})\bar{\bm{e}}\approx R_{i}^{1},\quad\textrm{and}\quad\nabla_{\bar{\bm{u}}}u^{+}_{i+1/2}(\bar{\bm{u}})\bar{\bm{e}}\approx\widetilde{R}_{i+1}^{1}.$ (54)

### v2

@@ 3 Nonlinear iteration scheme
(13) $\bm{r}(\bar{\bm{u}}):=\bm{b}-{\cal A}(\bar{\bm{u}})\equiv\begin{bmatrix}\bar{\bm{u}}^{0}\\ \bm{0}\\ \vdots\\ \bm{0}\end{bmatrix}-\begin{bmatrix}&I&\\ &-\Phi(\,\cdot\,)&I\\ &&\ddots&\ddots\\ &&&-\Phi(\,\cdot\,)&I\end{bmatrix}\begin{bmatrix}\bar{\bm{u}}^{0}\\ \bar{\bm{u}}^{1}\\ \vdots\\ \bar{\bm{u}}^{n_{t}-1}\end{bmatrix}=\bm{0}.$
Here $\bar{\bm{u}}=\big(\bar{\bm{u}}^{0},\bar{\bm{u}}^{1},\ldots,\bar{\bm{u}}^{n_{t}-1}\big)^{\top}\in\mathbb{R}^{n_{x}n_{t}}$ is created from concatenating the cell-averages across the time domain, ${\cal A}\colon\mathbb{R}^{n_{x}n_{t}}\to\mathbb{R}^{n_{x}n_{t}}$ is the nonlinear space-time discretization operator we seek to invert, and $\bm{b}\in\mathbb{R}^{n_{x}n_{t}}$ contains the initial data for the problem. For some approximate solution $\bar{\bm{u}}_{k}\approx\bar{\bm{u}}$, with iteration index $k$, the residual (13) is denoted as $\bm{r}_{k}:=\bm{r}(\bar{\bm{u}}_{k})$.
In this paper, we iteratively solve (13) using the linearly preconditioned residual correction scheme outlined in Algorithm 1. The remainder of this section discusses in greater detail the development of Algorithm 1; however, we first use it to contextualize the key contributions of this paper. In Algorithm 1, $P_{k}$ is a certain linearization of the space-time discretization $\cal A$} in (13), and the development of this linearization in Section 4 is the first key contribution of this work; specifically, $P_{k}$ is realized as a certain non-standard discretization of the linear conservation law (cons-lin). In Algorithm 1 linear systems of the form $P_{k}\bar{\bm{e}}_{k}^{\textrm{lin}}=\bm{r}_{k}$ are solved either directly with sequential time-stepping or approximately with MGRIT. We again emphasize that the use of MGRIT here is optional, and that any parallel-in-time solver which is effective on problems of the form (cons-lin) could be used in its place. In this work we focus on the MGRIT solution option in Algorithm 1 since we are concerned with parallel-in-time solvers. The development of an effective MGRIT solver for $P_{k}$ is the second key contribution of this work, and is the subject of Section 5. Our MGRIT solver is based on a novel, modified FV semi-Lagrangian coarse-grid discretization of (cons-lin) which is an extension of our modified finite-difference (FD) semi-Lagrangian discretizations in [14, 15, 16] for non-conservative linear hyperbolic PDEs. The efficacy of our solution methodology is demonstrated in Section 6 by way of numerical tests on challenging nonlinear problems containing shock and rarefaction waves.
@@
5:   ${\bar{\bm{e}}}_{k}^{\textrm{lin}}\leftarrow P_{k}^{-1}{\bm{r}}_{k}$ $\triangleright$ Direct solve (sequential in time)
@@
7:   ${\bar{\bm{e}}}_{k}^{\textrm{lin}}\leftarrow\textrm{MGRIT}(P_{k},{\bm{r}}_{k})$ $\triangleright$ Approx. solve (parallel in time)
@@
Since $\cal A$} in (13), and consequently $\nabla_{\bar{\bm{u}}}{\cal A}(\bar{\bm{u}})$, corresponds to a space-time discretization utilizing a one-step method in time, we impose that $P_{k}$ has the corresponding sparsity structure:
(15) $P_{k}=\begin{bmatrix}&I\\ &-\Phi^{\textrm{lin}}(\bar{\bm{u}}_{k}^{0})&I\\ &&\ddots&\ddots\\ &&&-\Phi^{\textrm{lin}}(\bar{\bm{u}}_{k}^{n_{t}-2})&I\end{bmatrix}\in\mathbb{R}^{n_{x}n_{t}\times n_{x}n_{t}}.$
@@
(16) $\Phi^{\textrm{lin}}(\bar{\bm{u}}^{n}_{k})=\nabla_{\bm{\alpha}}\widetilde{\Phi}(\bm{\alpha},\bm{\beta})\big|_{(\bm{\alpha},\bm{\beta})=(\bar{\bm{u}}^{n}_{k},\bar{\bm{u}}^{n}_{k})}.$
In the coming sections we detail specifically how $\widetilde{\Phi}$ is chosen (see e.g., (25)). Moving forward, to maintain a uniform and simple notation, we continue to refer to the linearized time-stepping operator $\Phi^{\textrm{lin}}(\bar{\bm{u}}^{n}_{k})$ in (16) as the “Jacobian of $\Phi$” and use the corresponding notation $\nabla\Phi$,” also when the Jacobian does not exist. In cases where it does not exist, “Jacobian of $\Phi$” and $\nabla\Phi$” are to be understood in the sense described above and as written in (16).
Finally, our numerical tests indicate that the linearized iteration (14) may benefit from the addition of nonlinear block relaxations of the type that would be done on (13) if the system were to be solved with FAS MGRIT. Hence we add Line 2 in Algorithm 1. To this end, let $m\in\mathbb{N}\setminus\{1\}$ induce a coarse-fine splitting (CF-splitting) of the time points such that every $m$th time point is a C-point, and all other time points are F-points. Then, nonlinear C- and F-relaxations update the current approximation $\bar{\bm{u}}_{k}$ such that the nonlinear residual is zero at C- and F-points, respectively [36]. A C-relaxation requires time-stepping to each C-point from its preceding F-point, and an F-relaxation requires time-stepping from each C-point to its succeeding m1m-1 F-points. In practice we only do an F-relaxation in Line 2, but note that any stronger relaxation ending in F is also possible.
@@ 4.1 Linearization I: Time integration
(17) $\big[\nabla_{\bar{\bm{u}}^{n}}\Phi(\bar{\bm{u}}^{n})\big]\bar{\bm{e}}^{n}=\big[\nabla_{\bar{\bm{u}}^{n}}F(\bar{\bm{u}}^{n})\big]\bar{\bm{e}}^{n}=\underbrace{\big[I+\delta t\nabla_{\bar{\bm{u}}^{n}}L(\bar{\bm{u}}^{n})\big]}_{=:F^{\textrm{lin}}(\bar{\bm{u}}^{n})}\bar{\bm{e}}^{n}.$
@@
(18) $\nabla_{\bar{\bm{u}}^{n}}\Phi(\bar{\bm{u}}^{n})\bar{\bm{e}}^{n}=\frac{1}{3}\bar{\bm{e}}^{n}+\frac{2}{3}\nabla_{\bar{\bm{u}}^{n}}F\big(\bar{\bm{u}}^{n,2}\big)\bar{\bm{e}}^{n}=\frac{1}{3}\bar{\bm{e}}^{n}+\frac{2}{3}F^{\textrm{lin}}\big(\bar{\bm{u}}^{n,2}\big)\underbrace{\nabla_{\bar{\bm{u}}^{n}}\bar{\bm{u}}^{n,2}\bar{\bm{e}}^{n}}_{=:\bar{\bm{e}}^{n,2}}.$
@@
(19) $\bar{\bm{e}}^{n,2}=\frac{3}{4}\bar{\bm{e}}^{n}+\frac{1}{4}F^{\textrm{lin}}\big(\bar{\bm{u}}^{n,1}\big)\underbrace{\nabla_{\bar{\bm{u}}^{n}}\bar{\bm{u}}^{n,1}\bar{\bm{e}}^{n}}_{=:\bar{\bm{e}}^{n,1}},\quad\textrm{and}\quad\bar{\bm{e}}^{n,1}=F^{\textrm{lin}}(\bar{\bm{u}}^{n})\bar{\bm{e}}^{n}.$
@@ 4.2 Linearization II: Numerical flux
(20) $-\nabla_{\bar{\bm{u}}}\big(L(\bar{\bm{u}})\big)_{i}\bar{\bm{e}}=\frac{\nabla_{\bar{\bm{u}}}\widehat{f}_{i+1/2}(\bar{\bm{u}})\bar{\bm{e}}-\nabla_{\bar{\bm{u}}}\widehat{f}_{i-1/2}(\bar{\bm{u}})\bar{\bm{e}}}{h}.$
@@
(21) $\nabla_{\bar{\bm{u}}}\widehat{f}_{i+1/2}^{\rm(global)}(\bar{\bm{u}})\bar{\bm{e}}=\nabla_{\bar{\bm{u}}}\widehat{f}^{\rm(global)}\big(u_{i+1/2}^{-}(\bar{\bm{u}}),u_{i+1/2}^{+}(\bar{\bm{u}})\big)\bar{\bm{e}},$
(22) $=\left(\frac{\partial\widehat{f}^{\rm(global)}}{\partial u_{i+1/2}^{-}}\nabla_{\bar{\bm{u}}}u_{i+1/2}^{-}(\bar{\bm{u}})+\frac{\partial\widehat{f}^{\rm(global)}}{\partial u_{i+1/2}^{+}}\nabla_{\bar{\bm{u}}}u_{i+1/2}^{+}(\bar{\bm{u}})\right)\bar{\bm{e}},$
(23) $\begin{split}&=\frac{1}{2}\left(f^{\prime}\big(u_{i+1/2}^{-}\big)+\nu^{\rm(global)}\right)\Big(\nabla_{\bar{\bm{u}}}u_{i+1/2}^{-}(\bar{\bm{u}})\bar{\bm{e}}\Big)\\ &\hskip 34.44434pt+\frac{1}{2}\left(f^{\prime}\big(u_{i+1/2}^{+}\big)-\nu^{\rm(global)}\right)\Big(\nabla_{\bar{\bm{u}}}u_{i+1/2}^{+}(\bar{\bm{u}})\bar{\bm{e}}\Big)\end{split},$
(24) $\begin{split}&=\widehat{f}^{\textrm{lin}}\Big(f^{\prime}\big(u_{i+1/2}^{-}\big),\nabla_{\bar{\bm{u}}}u_{i+1/2}^{-}(\bar{\bm{u}})\bar{\bm{e}},f^{\prime}\big(u_{i+1/2}^{+}\big),\nabla_{\bar{\bm{u}}}u_{i+1/2}^{+}(\bar{\bm{u}})\bar{\bm{e}},\nu^{\rm(global)}\Big).\end{split}$
@@
(25) $\begin{split}\widetilde{f}_{i+1/2}^{\rm(local)}(\bm{\alpha},\bm{\beta})&:=\frac{1}{2}\Big[f(u_{i+1/2}^{-}(\bm{\alpha}))+f(u_{i+1/2}^{+}(\bm{\alpha}))\Big]\\ &\hskip 34.44434pt+\frac{1}{2}\nu_{i+1/2}(\bm{\beta})\Big[u_{i+1/2}^{-}(\bm{\alpha})-u_{i+1/2}^{+}(\bm{\alpha})\Big]\end{split}.$
@@
(26) $\begin{split}&\nabla_{\bar{\bm{u}}}\widehat{f}_{i+1/2}^{\rm(local)}(\bar{\bm{u}})\bar{\bm{e}}=\nabla_{\bm{\alpha}}\widetilde{f}_{i+1/2}^{\rm(local)}(\bm{\alpha},\bm{\beta})|_{(\bm{\alpha},\bm{\beta})=(\bar{\bm{u}},\bar{\bm{u}})}\bar{\bm{e}}=\\ &\widehat{f}^{\textrm{lin}}\Big(f^{\prime}\big(u_{i+1/2}^{-}\big),\nabla_{\bar{\bm{u}}}u_{i+1/2}^{-}(\bar{\bm{u}})\bar{\bm{e}},f^{\prime}\big(u_{i+1/2}^{+}\big),\nabla_{\bar{\bm{u}}}u_{i+1/2}^{+}(\bar{\bm{u}})\bar{\bm{e}},\nu_{i+1/2}^{\rm(local)}(\bar{\bm{u}})\Big).\end{split}$
@@ 4.3 Linearization III: Reconstructions
(28) $\nabla_{\bar{\bm{u}}}u^{-}_{i+1/2}(\bar{\bm{u}})\bar{\bm{e}}$ $=\underbrace{\sum\limits_{\ell=0}^{k-1}b_{i}^{\ell}(\bar{\bm{u}})\big(R^{\ell}\bar{\bm{e}}\big)_{i}}_{=:R_{i}^{1}}+\underbrace{\sum\limits_{\ell=0}^{k-1}\big(R^{\ell}\bar{\bm{u}}\big)_{i}\big(\nabla_{\bar{\bm{u}}}\,b_{i}^{\ell}(\bar{\bm{u}})\bar{\bm{e}}\big)_{i}}_{=:R_{i}^{2}},$
(29) $\nabla_{\bar{\bm{u}}}u^{+}_{i+1/2}(\bar{\bm{u}})\bar{\bm{e}}$ $=\underbrace{\sum\limits_{\ell=0}^{k-1}\widetilde{b}_{i+1}^{\ell}(\bar{\bm{u}})\big(\widetilde{R}^{\ell}\bar{\bm{e}}\big)_{i+1}}_{=:\widetilde{R}_{i+1}^{1}}+\underbrace{\sum\limits_{\ell=0}^{k-1}\big(\widetilde{R}^{\ell}\bar{\bm{u}}\big)_{i+1}\big(\nabla_{\bar{\bm{u}}}\,\widetilde{b}_{i+1}^{\ell}(\bar{\bm{u}})\bar{\bm{e}}\big)_{i+1}}_{=:\widetilde{R}_{i+1}^{2}}.$
@@
(30) $k=1:\quad d^{0}=1,\quad\widetilde{d}^{0}=1,$
(31) $k=2:\quad\big(d^{0},d^{1}\big)=\Big(\frac{2}{3},\frac{1}{3}\Big),\quad\big(\widetilde{d}^{0},\widetilde{d}^{1}\big)=\Big(\frac{1}{3},\frac{2}{3}\Big).$
In this case $R_{i}^{1},\widetilde{R}_{i+1}^{1}$ are simply linear, (2k1)(2k-1)-order reconstructions of $e$. Moreover, the second terms in (28) and (29) vanish, $R_{i}^{2}=\widetilde{R}_{i+1}^{2}=0$, because the weights are independent of $\bar{\bm{u}}$. Recall that these optimal linear weights are used in a 1st-order accurate FV method. However, for higher-order discretizations, these weights are not suitable when the solution lacks regularity because they result in spurious oscillations.
@@
(32) $w^{\ell}_{i}(\bar{\bm{u}})$ $=\frac{\alpha_{i}^{\ell}(\bar{\bm{u}})}{\sum_{\ell=0}^{k-1}\alpha_{i}^{\ell}(\bar{\bm{u}})},\quad\alpha_{i}^{\ell}(\bar{\bm{u}})=\frac{d^{\ell}}{\big(\epsilon+\beta_{i}^{\ell}(\bar{\bm{u}})\big)^{2}},$
(33) $\widetilde{w}^{\ell}_{i+1}(\bar{\bm{u}})$ $=\frac{\widetilde{\alpha}_{i+1}^{\ell}(\bar{\bm{u}})}{\sum_{\ell=0}^{k-1}\widetilde{\alpha}_{i+1}^{\ell}(\bar{\bm{u}})},\quad\widetilde{\alpha}_{i+1}^{\ell}(\bar{\bm{u}})=\frac{\widetilde{d}^{\ell}}{\big(\epsilon+\beta_{i+1}^{\ell}(\bar{\bm{u}})\big)^{2}},$
[…]
(34) $\beta^{0}_{i}(\bar{\bm{u}})=(\bar{u}_{i+1}-\bar{u}_{i})^{2},\quad\beta^{1}_{i}(\bar{\bm{u}})=(\bar{u}_{i}-\bar{u}_{i-1})^{2}.$
@@
(35) $\nabla_{\bar{\bm{u}}}u^{\pm}_{i+1/2}(\bar{\bm{u}})\bar{\bm{e}}\approx\frac{u^{\pm}_{i+1/2}(\bar{\bm{u}}+\mu\bar{\bm{e}})-u^{\pm}_{i+1/2}(\bar{\bm{u}})}{\mu},$
@@
(36) $\nabla_{\bar{\bm{u}}}u^{-}_{i+1/2}(\bar{\bm{u}})\bar{\bm{e}}\approx R_{i}^{1},\quad\textrm{and}\quad\nabla_{\bar{\bm{u}}}u^{+}_{i+1/2}(\bar{\bm{u}})\bar{\bm{e}}\approx\widetilde{R}_{i+1}^{1}.$